OpenStudy (anonymous):

What is the equation of the line that passes through (1, 3) and (-2, -3)?

OpenStudy (anonymous):

first find the slope. -3 - 3 ------- -2 - 1 -6/-3 = 2 y = 2x + b plug in a point 3 = 2(1) + b so b = 0 so y = 2x

OpenStudy (anonymous):

y = 2x + 1

OpenStudy (anonymous):

3 = 2(1) + b so b = 0 <=== no b=1 so y = 2x => y =2x+1